Renu completed 2/3 part of her home work in 2 hours. How much part of her home work had she completed in 1¼ hours?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question it is given that,

Renu completed 2/3 part of her home work in 2 hours.Let us assume total part of homework be ‘P’.
Then,

2/3 of P = 2

2/3 × P = 2
By cross multiplication we get,

P = 2 × 3/2

P = 3 home work
So, part of her home work she had completed in 11⁄4 hours i.e. 5/4 hours Let us assume part of home work she had completed in 5/4 hours be = Q Then,

Q of 3 = 5/4

Q × 3 = 5/4
By cross multiplication we get,

Q = (5/4) × (1/3)

Q = 5/12
Therefore, Renu completed 5/12 part of her home work in 5/4 hours.


Thanks